OpenStudy (anonymous):

MY VERY LAST QUESTION FOR BIOLOGY PLEASE HELP ME ITS GOT ME STUCK!!! Which of the following best describes reproductive structures common to both gymnosperms and angiosperms? Seeds formed after combining pollen and egg cells during sexual reproduction Male pollen grains and female egg cells in the same reproductive organ Cones containing seeds that are eaten and carried away by animals Flowers designed to attract pollinators to assist in fertilization

OpenStudy (anonymous):

i think it is A...

OpenStudy (abhisar):

D and A are false !

OpenStudy (abhisar):

no wait !

OpenStudy (anonymous):

so if its not a i was also thinking B.

OpenStudy (anonymous):

ok

OpenStudy (abhisar):

No it's A !

OpenStudy (abhisar):

Both gymno and angio produce seeds but angio produces flowers and fruits which is absent in gymno.

OpenStudy (anonymous):

yep

OpenStudy (abhisar):

Well done ! @ilovehim121511

OpenStudy (anonymous):

okay lol i thought so you had me second guessing myself.. (:

OpenStudy (anonymous):

IM SO NERVOUS! I HOPE I PASSED!!

OpenStudy (abhisar):

ouchhhh!! is it a test ?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

no i already did everything this was my last assignment. but im done with BIOLOGY NOWW ((:

OpenStudy (abhisar):

Yeah..i'll not charge you cz u were ryt ! But that is wrong...to take external help while test and also it is against the code of conduct. Tests are for ur self assessment. All the Best ! \(\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \Large\cdot \quad \cdot\\ \cdot\\ \huge \smile \end{matrix} \right)}}\normalsize \\ \;/\quad \;\;\quad \backslash\)

OpenStudy (anonymous):

this one wasnt a test lol. it was my assignment that i hadent done from a while ago.

OpenStudy (abhisar):

ok \(\color{green}{\huge\ddot\smile}\)

OpenStudy (anonymous):

Lol. This is indeed a test. @ilovehim121511 @Abhisar